### v1

# LLaMIPa: An Incremental Discourse Parser

Abstract
This paper provides the first discourse parsing experiments with a large language model (LLM) finetuned on corpora annotated in the style of SDRT (Segmented Discourse Representation Theory Asher,, 1993; Asher and Lascarides,, 2003). The result is a discourse parser, LLaMIPa (LLaMA Incremental Parser), which is able to more fully exploit discourse context, leading to substantial performance gains over approaches that use encoder-only models to provide local, context-sensitive representations of discourse units. Furthermore, it is able to process discourse data incrementally, which is essential for the eventual use of discourse information in downstream tasks.
1 Introduction
Understanding a conversation requires more than just grasping the contents of its utterances. One must also understand how different utterances relate to one another: whether an utterance answers a question or serves to explain or correct something else that was said, for example. Discourse parsing, the task of predicting such relations, often represents the discourse structure of texts or conversations as graphs in which elementary discourse units or EDUs (roughly clause-level contents) are represented by nodes, and semantic relations between clauses are represented as labeled edges or separate nodes.
Discourse parsing remains a challenge in NLP due to three factors: (i) the complexity of discourse graphs, (ii) the frequent lack of surface cues provided by EDUs, which forces parsers to rely on deep, semantic information, (iii) the context dependent nature of the inference required to predict the graph elements.
Recently, improved discourse parsers have incorporated neural models and encoder-only pretrained language models like BERT Devlin et al., (2018), which offer richer representations of linguistic content, and can address factor (ii). But given their limited input size, they are confined to look at one small part of a discourse at a time and so cannot really address factor (iii).
Generative large language models (LLMs) can substantially improve discourse parsing. They not only encode large amounts of lexical and conceptual information, addressing problem (ii) but their increased context windows permits an incremental parsing strategy to address problem (iii). Such a strategy makes sense cognitively and linguistically, as prior discourse structure clearly has an effect on inferences about which subsequent discourse units are attached (the link task) and how they are attached—i.e. what semantic relations link them (the link+relation task) Asher and Lascarides, (2003). Further, an incremental approach naturally applies to downstream applications like conversational agent, in which the information about the previously inferred discourse structure must be updated and available in real time.
We develop an incremental parsing model LLaMIPa (LLaMa Incremental Parser), by fine-tuning LLaMA3, that uses previously inferred discourse structure to simultaneously predict the link and the link+relation tasks between new discourse units. We show LLaMIPa substantially outperforms the state of the art on both tasks.
We demonstrate LLaMIPa’s performance using three datasets annotated in the style of SDRT (Segmented Discourse Representation Theory Asher,, 1993; Asher and Lascarides,, 2003). Our central dataset is the Minecraft Structured Dialogue Corpus (MSDC; Thompson et al.,, 2024), a corpus of situated, chat-based, task-oriented dialogues, but we show the generality of our approach by separately fine-tuning on STAC, a corpus of situated multi-party chats from an online game Asher et al., (2016). We further test (STAC-based) LLaMIPa on an out of domain dataset, Molweni Li et al., (2020).
Not only does LLaMIPa outperform prior models, but it handles particularly complex structural features of our situated, multiparty datasets, addressing problem (i). In particular, LLaMIPa can identify cases in which a single EDU has two parent EDUs—a situation that arises naturally in both multiparty and situated dialogue where a speaker can respond to multiple channels (conversational or multimodal) at once. Most current SDRT based parsers cannot handle such structures Shi and Huang, (2019); Liu and Chen, (2021); Wang et al., (2021); Chi and Rudnicky, (2022).
In what follows, Section 2 provides background on discourse structure. Section 3 presents related work, while Section 4 details the multiparty corpora we use for training and testing. Sections 5 and 6 describe our model and results. Section 7 describes ablation experiments, and then we conclude in Section 8.
2 Background
Discourse representations for a text or conversation often take the form of tree or graph structures that are built recursively from EDUs, where EDUs are clauses or subclausal units that serve as minimal, linguistic constituents in discourse relations Marcu, (1999). Such relations, such as Explanation, Question-Answer Pair or Acknowledgement, are represented as linking EDUs in the discourse graph.
There are two main theories that have investigated complete discourse structures for texts: Rhetorical Structure Theory (RST; Mann and Thompson,, 1987) and SDRT (Asher,, 1993; Asher and Lascarides,, 2003). RST builds discourse trees over EDUs with the semantic relation labels on the head nodes dominating subtrees.
SDRT provides a dependency style, weakly connected directed acyclic graph (DAG) structure in which typed relations are represented as labeled arcs in the graph, and the terms of the relations are EDUs or elementary event units (EEUs). EEUs represent salient aspects of the nonlinguistic context in which a conversation takes place, and may be connected to the discourse DAG. SDRT also allows subgraphs, or complex discourse units (CDUs) to act as terms of semantic relations.
Discourse parsing requires addressing three tasks. The segmentation task involves decomposing a text and (descriptions of) actions into EDUs and EEUs. The link task involves identifying pairs of EDUs and/or EEUs that are semantically related via discourse relations—that is, the arcs in the discourse graph. The link+relation task involves identifying the labels for each attached pair.
3 Related work
Traditionally with SDRT style parsers, gold segmentation is assumed Muller et al., (2012); Afantenos et al., (2015); Perret et al., (2016); Shi and Huang, (2019); Liu and Chen, (2021); Chi and Rudnicky, (2022), and parsing follows a pipeline architecture in which discourse units are first linked and then labels are subsequently predicted for these links. While we assume gold segmentation, our proposed architecture combines the attachment and labeling tasks.
Several strategies have evolved to take account of the complex information required for discourse parsing within a dependency based paradigm like that proposed by SDRT or DLTag (Forbes et al.,, 2003). Research has shown that semantic features of EDUs are crucial for solving the link+relation task Mann and Thompson, (1987); Asher and Lascarides, (2003). Older approaches to multiparty conversation Afantenos et al., (2015); Perret et al., (2016) thus used manual features about EDU pairs and simple ML models to build a local model to predict link and link+relation label in a pipeline fashion. A local model uses information just about pairs of EDUs that are to be attached; it does not exploit information about the discourse structure around that pair. Afantenos et al., (2015); Perret et al., (2016) also added decoding mechanisms with post hoc constraints on the output structures, following Muller et al., (2012). Although exploiting little to no information about other components of the discourse structure, such a model can still be competitive Bennis et al., (2023). We do a detailed comparison between an encoder based local parsing model and LLaMIPa.
Shi and Huang, (2019) were the first to obtain significant discourse parsing results using a deep learning architecture on STAC. They attempted to capture incremental and contextual effects in their model by training a supplemental Structured Encoder to incrementally update attachment paths (sequences of parent-child EDUs). However, Wang et al., (2021) showed that the model obtains similar scores with or without this encoder.
Liu and Chen, (2021) develop an approach that uses a pre-trained RoBERTa model Liu et al., (2019) to provide embeddings for EDUS. In addition, they use a bi-GRU cell Cho et al., (2014) to capture contextual information; but this requires them to limit the size of the input. Their model uses two linear layers for link and relation prediction.
Wang et al., (2021) provide the Structure Self-Aware Graph Neural Network (SSA-GNN), which is a complex GNN-based architecture and model that uses both EDU and edge embeddings. The model uses a Hierachichal GRU gate for contextual EDU representations. They then apply the SSA-GNN to capture implicit structural information between EDUs, using a Structure-Aware Scaled Dot-Product Attention Zhu et al., (2019); Wang et al., (2020) to update edge and EDU representations. They further train a teacher network to supplement the standard classification loss with an auxiliary loss. Chi and Rudnicky, (2022) developed a more complex model based on Shi and Huang, (2019).
None of the models by Shi and Huang, (2019); Liu and Chen, (2021); Guz et al., (2020); Wang et al., (2021); Chi and Rudnicky, (2022) predict multiple parents for an EDU (or EEU). The only previously developed parser that predicts multiple parents during attachment is Bennis et al., (2023); Thompson et al., (2024)’s model BERTLine. BERTLine uses BERT embeddings for EDUs and EEUs, along with a linear layer for features like speaker change and distance information, to build a local model and predicts link and link+relation in pipeline fashion. However the relation labeling uses a multitask architecture to capture informational dependencies between link and relation labeling decisions. Bennis et al., (2023) show that their model is competitive with approaches that try to build in more contextual knowledge Liu and Chen, (2021); Wang et al., (2021). Thompson et al., (2024) uses BERTLine on the MSDC, which makes it a natural baseline for our work on situated, multiparty conversation.
Recent work has investigated using LLMs to predict discourse structure, by finetuning LLaMA2 7b, 13b, and 70b for RST-style parsing Maekawa et al., (2024). The reported results on RST corpora are not directly comparable to ours, but they are significantly lower numerically. Chan et al., (2024); Yung et al., (2024); Fan et al., (2024); Henning and Stede, (2024) have investigated discourse parsing through prompting with large GPT-style models to produce PDTB (Prasad et al.,, 2008) or other discourse annotations. The results are significantly worse than dedicated discourse parsers; testing with 3-shot learning on a version of the STAC corpus (STAC-L; see Section 4), Chan et al., (2024) reports an F1 score of 21.3 for link prediction and 7.4 for relation prediction by ChatGPT.
4 Data sets
MSDC. The MSDC provides full discourse structures in the style of SDRT to dialogues in the Minecraft Dialogue Corpus (MDC; Narayan-Chen et al.,, 2019). In the MDC, dialogues involve a builder, whose task is to build block formations in a simplified Minecraft world, and an architect, who guides the builder by giving instructions through chat. The MSDC annotates discourse relations between both chat moves (EDUs) and linguistic descriptions of the builder’s pick and place actions (EEUs), provided by the MDC logs.
The MSDC’s multimodal nature presents many opportunities to simultaneously respond to a conversational move and a building move, and indeed, the MSDC has numerous discourse units with multiple parents (MPDUs). Figure 1 shows an excerpt from the MSDC with two examples. For instance, after placing two blocks, the builder asks “like that?”. The architect’s reply, “not quite,” not only serves to answer the builder’s question but simultaneously to correct the builder’s last moves. The following two elaborating discourse moves fill out the full content of the Correction.
Figure 1 shows another important aspect of the MSDC, namely that builders tend to perform sequences of pick and place moves uninterrupted by chat moves. These action sequences are grouped together as CDUs in the MSDC, indicated by dashed boxes in Figure 1. Bennis et al., (2023), Thompson et al., (2024) merged these highly repetitive sequences of nonlinguistic moves into a single EEU to prevent them from drowning out the more complex and interesting linguistic interactions in the corpus. We used the MSDC with this modification.
| DU-type | STAC-L | STAC-Sit | MSDC |
| Train/Test | Train/Test | Train/Test | |
| EDU | 11398/1154 | 11707/1163 | 17135/5402 |
| EEU | 0/0 | 11566/1156 | 4687/1473 |
| MDPU | 838//86 | 999/96 | 4789/1476 |
| MPDU-3 | 94/9 | 95/6 | 85/35 |
| MPDU>3 | 11/0 | 7/0 | 1/3 |
STAC. STAC111https://https://www.irit.fr/STAC/corpus.html is a corpus of multiparty chats from the game Settlers of Catan, in which players trade or otherwise acquire resources in order to build roads and settlements and thereby score victory points. STAC has two versions: STAC-L, in which only relations between chat moves (EDUs) are annotated in the absence of the context provided by nonlinguistic events, and a situated version, STAC-Sit, in which both chat moves and (descriptions of) game moves (EEUs) are annotated in the presence of an integrated linguistic and nonlinguistic context.
STAC contains MPDUs due to both multimodal interactions and multiparty discourse, in which one player can simultneously respond to multiple players at once. Relevant statistics for STAC and the MSDC are found in Table 1.
STAC-Sit, like the MSDC, also contains sequences of game-only moves. Bennis et al., (2023) compressed such sequences into individual EEUs and we follow this choice. Unlike the MSDC, however, STAC contains numerous CDUs for chat moves and for chat/game interactions (as opposed to CDUs of action sequences only). Following Bennis et al., (2023) and most of the literature on SDRT parsing, we did not treat CDUs directly but used the “flattening" procedure outlined in Muller et al., (2012) that converts a structure with CDUs into one with only EDUs and EEUs. A final simplication of the original STAC corpus involved removing EEUs that neither interact with EDUs n or figure in EEU sequences that interact with EDUs. This choice, like the one to compress EEU sequences, was made to avoid drowning out the more complex and interesting data involving linguistic EDUs and their interactions with EEUs.
Molweni. The Molweni corpus contains SDRT graphs for multiparty interactions in the Ubuntu Chat Corpus Lowe et al., (2015) a collection of exchanges from an online forum for discussing issues related to Ubuntu. Unlike the MSDC and STAC, Molweni does not contain CDUs, EEUs, or MPDUs. We used the version of Molweni Test re-annotated by Li et al., (2024).
5 LLaMIPa: An Incremental Parser
LLaMIPa leverages the sequence to sequence architecture and the enhanced contextual reasoning capacities of generative LLMs by introducing an incremental discourse parsing strategy. As it generates a discourse graph for a conversation , the LLM should base its link and relation predictions for an EDU in on the graph it has already constructed for EDUs preceding .
Formally, let be a conversation consisting of a sequence of EDUs or EEUs. Our LLM takes as input a set of of discourse and (possibly) event units and a discourse graph . At step 0, is , the head EDU, and , the null discourse graph. After that, is a set , where 1 and where is the window size, set to ; and contains links for . At each step, the model predicts a graph that extends by integrating link+relation predictions for all of the EDUs in the next dialogue turn or, if the next move is an action sequence, the EEU representing this sequence. This process repeats until LLaMIPa arrives at the final unit of , effectively creating a “sliding context window” that will eventually slide over an entire dialogue, permitting the model to potentially exploit a large amount of contextual information.
We represent relation instances as EDU index pairs connected by a relation type, e.g. RES(0,1), for the instance of a Result relation holding between EDUs 0 and 1. During training, each subsequent instance in a dialogue adds to the context the previous turn EDUs and EEUs, but we take the gold, rather than predicted, discourse structure between those units to make our input graph . is specified as a list of the relations between these units in the same format as the predicted relations.
We used a sliding context widow of 15 EDUs to optimize the number of training samples while capturing the most number of discourse links that have statistical support. As discourse links follow an inverse logarithmic curve Afantenos et al., (2015), there is little statistical support for links of distance longer than 15 in the SDRT corpora.
Using our corpora formatted for incremental relation prediction, we finetuned LLaMA-2 Touvron et al., (2023), and LLaMA-3222https://llama.meta.com/llama3/ to produce LLaMIPa2 and LLaMIPa3. In order to do this without training the full model weights, we use QLoRA Dettmers et al., (2024) to train for 3 epochs. The model parameter details are in Table 6 in Appendix A. The finetuned models will be made publicly available.
Once LLaMIPa is finetuned on the training corpus, the generation procedure uses predicted discourse structure instead of the gold structures during inference (See Figure 2). Proceeding turn by turn, LLaMIPa produces a complete discourse structure for a dialogue in which both links and relations are predicted. The LLaMIPa3 model and generation code is available here333https://huggingface.co/Kqte/LLaMIPa.
6 Results
We first ran our finetuning experiments on the MSDC corpus, and assessed LLaMIPa’s performance on link and relation type prediction when we include the gold dialogue structure in each sample context. We then tested LLaMIPa by using the previously predicted relations in the context instead of the gold ones, as in the generation step described above, in order to have our incremental parser build a predicted structure from beginning to end. Table 2 shows that LLaMIPa3 significantly outperforms BERTLine and outperforms LLaMIPa2 both using gold structure (LLaMIPa3+g vs. LLaMIPa2+g) and predicted structure (LLaMIPa3+p vs. LLaMIPa2+p). Below we use the better performing LLaMIPa3+p for other evaluations.
To test the generalizability of the model, we also finetuned LLaMIPa3 on STAC-Sit and compared it to BERTLine on MSDC and STAC-Sit. For both link prediction and link relation prediction, LLaMIPa shows a gain of almost 10 F1 points or 11% on link and 10.5 F1 points or 15% on link relation labeling. LLaMIPa posted even more impressive gains for predictions of certain relations. When we look at F1 breakdown by relation type in Table 5, the largest relative increase comes from improved predictions for the Narration and Correction relation types, which Thompson et al., (2024) cite as the most difficult to predict given their particular characteristics.
We then compared our model to a broader range of SDRT discourse parsers on the STAC linguistic version (STAC-L), as previous work focused on this version of the corpus. Table 4 shows that LLaMIPa3 finetuned on STAC-L and evaluated on STAC-L Test scored significantly higher on link prediction and got a statistically significantly higher score on link + relation prediction over the state-of-the-art parser Chi and Rudnicky, (2022).
Nevertheless, LLaMIPa3 has substantially lower performance on STAC-L (linguistic) than on either of the situated corpora. This mirrors human efforts at annotation. Asher et al., (2016, 2021) report that annotating the STAC linguistic data set was more difficult; annotators felt that information needed to determine the discourse structure was missing.
We also investigated LLaMIPa3 for transfer learning, finetuning on STAC-L and testing on the Molweni Cleaned Test corpus of Li et al., (2024). While superior to other models, in comparison to Li et al., (2024), LLaMIPa3 fared worse on link prediction but was far superior with respect to link + relation prediction.
Overall, LLaMIPa has several advantages. It simplifies previous attach-then-type pipeline methods by predicting link and relation labels simultaneously. Previous research has given theoretical arguments that co-predicting attachment and relation labels should give better results Asher and Lascarides, (2003), as well as empirical evidence for this claim Bennis et al., (2023). Our model also outperforms prior models on all relevant corpora and pretty much all tasks we studied. We hypothesize that this performance is due to the richer contextual information supplied to LLaMIPa, a hypothesis we now explore further.
| MSDC | STAC Situated | |||
| Link | Link+Rel | Link | Link+Rel | |
| LLaMIPa2+g | 0.8561 | 0.7664 | - | - |
| LLaMIPa2+p | 0.8579 | 0.7570 | - | - |
| LLaMIPa3+g | 0.9004 | 0.8154 | - | - |
| LLaMIPa3+p | 0.8830 | 0.7951 | 0.8612 | 0.7796 |
| BERTLine | 0.7870 | 0.6901 | 0.7667 | 0.6788 |
| Distance | 2 | 3 | 4 | 5 | 6 | 7 | 8 | 9 | 10 | 11 | 12 | 13 | 14 | 15 |
| LLaMIPa-3 | 0.90 | 0.88 | 0.82 | 0.76 | 0.82 | 0.80 | 0.79 | 0.71 | 0.90 | 0.76 | 0.50 | 0.40 | 1.0 | n/a |
| BERTLine | 0.70 | 0.61 | 0.43 | 0.32 | 0.38 | 0.02 | 0.21 | 0.28 | 0.08 | n/a | n/a | n/a | n/a | n/a |
| STAC-L | STAC- L/Molweni-clean | |||
| Link | Link+Rel | Link | Link+Rel | |
| Perret et al., (2016) | 0.69 | 0.531 | - | - |
| Shi and Huang, (2019) | 0.719 | 0.536 | - | - |
| Liu and Chen, (2021) | 0.731 | 0.571 | - | - |
| Wang et al., (2021) | 0.734 | 0.573 | - | - |
| Fan et al., (2022) | 0.736 | 0.574 | - | - |
| Chi and Rudnicky, (2022) | 0.744 | 0.596 | 0.645 | 0.380 |
| Li et al., (2024) | 0.593 | 0.386 | 0.756 | 0.312 |
| Bennis et al., (2023) (BERTLine) | 0.730 | 0.562 | - | - |
| Fan et al., (2024) (ChatGPT) | 0.599 | 0.252 | - | - |
| Chan et al., (2024) (ChatGPT) | 0.213 | 0.074 | - | - |
| LLaMIPa-3+p | 0.775 | 0.607 | 0.712 | 0.405 |
| BERTLine | LLaMA3-Local | LLaMIPa3+p | LLaMIPa3+g | LLaMIPa-rand | No Structure | |
| Result | 0.85 | 0.86 | 0.90 | 0.91 | 0.83 | 0.72 |
| Acknowledgement | 0.81 | 0.85 | 0.86 | 0.86 | 0.81 | 0.72 |
| Narration | 0.50 [0.73] | 0.54 | 0.82 | 0.91 | 0.28 | 0.18 |
| Elaboration | 0.75 | 0.77 | 0.77 | 0.77 | 0.75 | 0.68 |
| Correction | 0.31 | 0.64 | 0.73 | 0.80 | 0.49 | 0.52 |
| Continuation | 0.44 | 0.47 | 0.49 | 0.50 | 0.44 | 0.33 |
| Question-answer Pair | 0.76 | 0.80 | 0.82 | 0.83 | 0.42 | 0.14 |
| Comment | 0.50 | 0.57 | 0.60 | 0.61 | 0.57 | 0.54 |
| Confirmation-Question | 0.86 | 0.89 | 0.93 | 0.93 | 0.91 | 0.89 |
| Clarification-Question | 0.61 | 0.66 | 0.73 | 0.73 | 0.70 | 0.41 |
| Contrast | 0.80 | 0.79 | 0.75 | 0.74 | 0.72 | 0.68 |
| Question-Elaboration | 0.39 | 0.36 | 0.30 | 0.36 | 0.33 | 0.33 |
| Alternation | 0.88 | 0.88 | 0.83 | 0.88 | 0.90 | 0.96 |
| Explanation | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 |
| Conditional | 0.58 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 |
| Sequence | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.05 |
| Link+Rel F1 | 0.69 [0.71] | 0.73 | 0.80 | 0.81 | 0.65 | 0.56 |
| Link F1 | 0.78 [0.80] | 0.82 | 0.88 | 0.90 | 0.77 | 0.72 |
7 The role of context
A particularly remarkable demonstration of LLaMIPa’s discourse parsing prowess is its success in predicting Narration instances in the MSDC. Narrations are frequently quite long (with a significant proportion above distance 10), and so difficult to capture by local parsers like BERTLine Thompson et al., (2024). Table 3 shows Narration F1 at distances between 3 and 14 EDUs. Like other local parsers, BERTLine requires a distance cutoff that balances ignoring longer relations with mitigating the class imbalance caused by numerous unlinked candidates; for SDRT corpora, this is distance 10. As explained in Section 5, LLaMIPa allows for training up to distance 15, and impressively, it continues to make (somewhat steadily good) predictions up to distance 15.
LLaMIPa also has very strong performance in predicting instances of the Correction relation, which also has a significant number of long distance instances in the MSDC. Table 5 shows that its performance is more than double BERTLine’s on this relation; its F1 score for Correction is over 42 F1 points more than BERTLine’s.
We hypothesize that LLaMIPa has a high F1 on Narration and Correction because it has access to rich semantic information about the discourse context. It can use the meaning of previous discourse moves as well as an explicit representation of their discourse structure. This hypothesis coheres with our understanding of how Narration functions in the MSDC. Narrations connect high-level instructions between which participants clarify, question and elaborate on those instructions, as well as correct missteps (see Figure 1). While there is some statistical similarity in their surface forms, the important cues for isolating a new instruction, and hence a Narration attachment, involve certain patterns of Result and Acknowledgement links in the prior context, as well as the presence of (nonlinguistic) block placements. This pattern is regular enough that one can find some Narrations algorithmically as in Thompson et al., (2024)’s “Second Pass”: taking predictions from a first pass of the local model, Thompson et al., then apply an algorithmic approach to predict long distance Narration links from the local model’s predictions. We also tried this second pass for LLaMIPa, however there was no improvement on F1 for Narration scores, which shows that LLaMIPa had already learned this rule during training.
Besides its performance on Narration and Correction, LLaMIPa offers significant improvements over BERTLine for many other relation types, even those whose instances are typically not very long distance. Table 5 shows LLaMIPa has significantly higher scores for Result, Acknowledgment, Elaboration, Question-answer Pair, Clarification Question and Comment. We hypothesize that predictions on these too profit from complex contextual information.
LLaMIPa produces poorer results for relations such as Contrast, Alternation and Question-Elaboration. This may be because instances of these relations can be inferred reliably just from information given by the related EDUs or EEUs.
First ablation: embeddings vs. context. While the previous discussion suggests that LLaMA’s larger context window is responsible for its better performance, we wanted to check this hypothesis further. So we finetuned LLaMA3-8b over pairs of EDUs as input to predict links and relation types between pairs of EDUs. This produces a local model roughly comparable to BERTLine’s, though BERTLine’s architecture is different. Table 5 shows that, compared to BERTLine, LLaMA3-Local gave stronger performance without using the prior discourse context for link and relation prediction. However, when compared to LLaMIPa3, LLaMA3-Local has significantly lower link and link+relation labeling predictions overall. LLaMIPa3’s relation predictions for many relations are also significantly better—indeed sometimes vastly superior as in the F1 for Narration. This supports our conclusion that although LLaMA’s superior embeddings play a role, many relations, and especially relations like Narration, require more contextual information like that available to LLaMIPa.
Further ablations: removing structure. Given the result of our first ablation experiment, we wanted to investigate further whether LLaMIPa was using the discourse structure for its predictions, or was solely relying on the longer input sequence of EDUs and EEUs.
Our set up for LLaMIPa makes the assumption that it is the prior discourse structure that is relevant. In Table 5, there is a small but consistent drop between gold and predicted contexts in F1 performance for all relation types. This supports the conclusion that LLaMIPa takes the discourse structure of its context into account and that the quality of that structure affects prediction.
To provide more support for this conclusion, we performed another ablation in which we perturbed the structure at each increment by randomizing both the relation type and EDU index for each relation. We then tested LLaMIPa (finetuned in the usual way) on this input. LLaMIPa-rand showed a significant drop in performance from LLaMA3-Local for link and across almost all relation types (with the exception of Contrast and Question Elaboration). The drop for Narration and Correct is sharp, as well as for Question-answer Pair. We then performed a variant ablation, in which we removed all structure from LLaMIPa’s input context, leaving only a sequence of EDUs and EEUs. The performance of LLaMIPa on this data set (in the column “No Structure” of Table 5) shows another significant drop both in link and in link+relation predictions. The drop in relation labeling in addition holds for almost all relations, and a dramatic drop, especially for Narration, Correction, Clarification Question and Question-Answer Pair. This provides strong evidence that LLaMIPa needs to the use discourse structure and not just the EDU/EEU input sequence of the context to optimize its predictions for links and relation labels and that discourse structure is crucial for predicting some relations.
8 Conclusions
In this paper, we presented LLaMIPa3, an incremental SDRT style discourse parser that uses its prior discourse structure predictions to predict links and relation labels for new discourse units. We finetuned both LLaMa2-13b and LLaMA3-8b models to produce LLaMIPa2 and LLaMIPa3. We showed that LLaMIPa3 performs better than other parsers that use SDRT data sets. It excels in particular on data sets that feature conversation that contains both linguistic moves and nonlinguistic contexts that help interpret them.
We also compared LLaMIPa3 with a local model discourse parser based on LLaMA3-8b as well as on variants of LLaMIPa3 that had differing levels of access to the discourse context. LLaMIPa3 had significantly higher F1 scores for attachment and relation labeling on the data sets we used for evaluation than all of these.
The differential performance of LLaMIPa3 on STAC-L versus STAC-Sit, which reflects impressions of human annotators, suggests that nonlinguistic contexts are important for stable interpretation of situated conversation. We humans typically invent a nonlinguistic context for a conversation when one isn’t given. But such an invention is subjective and may contribute to variations in understanding and prediction. We suspect something like this might be true for LLMs too. In future work, we hope to use LLaMIPa3 to help conversational robots to improve their performance in interactive tasks.
Limitations
We did not exhaustively test all models on the situated data sets. Given the number of MPDUs especially in the MSDC, however, we felt that their performance would not compare to LLaMIPa’s. We plan to do more extensive testing in the future to compare models.
LLaMIPa, like previous parsers, presumes that the edus/eeus are already segmented. However extant work on automatic segmentation suggests that this task is relatively tractable on diverse data sets Muller et al., (2019) including even corpora of spoken conversation (Gravellier et al.,, 2021; Prevot et al.,, 2023)
Though ablations show that LLaMIPa uses discourse structure, it remains unclear how it leverages that structure to predict relations like Narration, Corrections, and QAP.
Like all supervised models, LLaMIPa is limited to making predictions on data sets that are similar to its training data. Given the heavy annotation effort required to build data sets annotated for discourse structure, this arguably puts a bottleneck on the generalizability of our approach to data sets in languages other than English and for conversation styles, such as multiparty spoken conversation, that are significantly different from the chat-based dialogues in the MSDC and STAC. Future work will explore efficient ways of harnessing extant data sources through, for instance, automatic translation, data augmentation and or/heuristics, to facilitate the transfer to other languages and domains.
Ethics Statement
Our work here has been to improve the capacities of AI models to understand the semantic structure of discourse and conversation. We see no direct ethical concerns that arise from this work, though we hope that a better grasp of the semantic structure of conversation may enable models to eventually produce less pragmatically grounded or conversational hallucinations.
Appendix A Appendix: Model Parameters
| GPUs | |
| 4 NVIDIA Volta V100 | |
| Hyperparameters | |
| Training epochs | 3 |
| batch size | 4 |
| optimizer | Adam |
| learning rate | 2e-4 |
| learning rate scheduler | linear warm-up and |
| cosine annealing | |
| warm-up ratio | 0.03 |
| gradient clipping | 0.3 |
| lora r | 64 |
| lora (alpha) | 16 |
| lora dropout ratio | 0.1 |
| lora target modules | Only Attention Blocks |
| (q_proj, v_proj) | |
| quantization for LLaMA3 | 4-bit NormalFloat |
Table 6 gives the hyperparameters used for finetuning LLaMIPa-2 and LLaMIPa-3 along with the computing resources. We adapted the finetuning code from the following repository444https://github.com/mlabonne/llm-course/blob/main/Fine_tune_Llama_2_in_Google_Colab.ipynb. We provide LLaMIPa3 trained on MSDC as well as the MSDC test data here555https://huggingface.co/Kqte/LLaMIPa
Appendix B Appendix: STAC Linguistic data
We constructed STAC-L, our STAC linguistic corpus, from the STAC data available on the STAC website (https://https://www.irit.fr/STAC/corpus.html) and adopted standard procedures for flattening CDUs. However, we found that our data set did not quite match that of Bennis et al., (2023) in terms of numbers of EDUs. Shi and Huang, (2019) also report a different number of EDUs from what we have and from what Bennis et al., (2023) report. When we finetuned and tested LLaMIPa on the STAC linguistic only data available in the BERTLine repository 666https://github.com/zineb198/LineBert, we had lower scores than what we report in Table 4. LLaMIPa-3+p had an F1 for link of 75.4 and a link+relation F1 of 57.7 on this version of STAC linguistic. On the transfer learning from STAC linguistic to Molweni Cleaned Test, LLaMIPa-3+p had an F1 of 68.5 for link and 38.7 for link+relation.

### v2

@@
 This paper provides the first discourse parsing experiments with a large language model (LLM) finetuned on corpora annotated in the style of SDRT (Segmented Discourse Representation Theory Asher,, 1993; Asher and Lascarides,, 2003).
-The result is a discourse parser, LLaMIPa (LLaMA Incremental Parser), which is able to more fully exploit discourse context, leading to substantial performance gains over approaches that use encoder-only models to provide local, context-sensitive representations of discourse units.
-Furthermore, it is able to process discourse data incrementally, which is essential for the eventual use of discourse information in downstream tasks.
+The result is a discourse parser, Llamipa (Llama Incremental Parser), that leverages discourse context, leading to substantial performance gains over approaches that use encoder-only models to provide local, context-sensitive representations of discourse units.
+Furthermore, it can process discourse data incrementally, which is essential for the eventual use of discourse information in downstream tasks.
 1 Introduction
 Understanding a conversation requires more than just grasping the contents of its utterances.
-One must also understand how different utterances relate to one another: whether an utterance answers a question or serves to explain or correct something else that was said, for example.
-Discourse parsing, the task of predicting such relations, often represents the discourse structure of texts or conversations as graphs in which elementary discourse units or EDUs (roughly clause-level contents) are represented by nodes, and semantic relations between clauses are represented as labeled edges or separate nodes.
-Discourse parsing remains a challenge in NLP due to three factors: (i) the complexity of discourse graphs, (ii) the frequent lack of surface cues provided by EDUs, which forces parsers to rely on deep, semantic information, (iii) the context dependent nature of the inference required to predict the graph elements.
-Recently, improved discourse parsers have incorporated neural models and encoder-only pretrained language models like BERT Devlin et al., (2018), which offer richer representations of linguistic content, and can address factor (ii).
-But given their limited input size, they are confined to look at one small part of a discourse at a time and so cannot really address factor (iii).
+One must also understand how different utterances relate to one another, for example whether an utterance answers a question or serves to explain or correct something previously said.
+Discourse parsing, the task of predicting such relations, often represents the discourse structure of texts or conversations as graphs in which elementary discourse units or EDUs (roughly, clause-level contents) are represented by nodes, and semantic relations between clauses are represented as labeled edges.
+Discourse parsing remains a challenge in NLP due to three factors: (i) the complexity of discourse graphs, (ii) the frequent lack of surface cues provided by EDUs, which forces parsers to rely on deep, semantic information, (iii) the context dependent nature of the inferences required to predict the graph elements.
+Recently, improved discourse parsers have incorporated neural models and encoder-only pretrained language models like BERT Devlin et al., (2019); Bennis et al., (2023) which offer richer representations of linguistic content, and can address factor (ii).
+But given their limited input size, they are confined to look at one small part of a discourse at a time, and so cannot really address factor (iii).
 Generative large language models (LLMs) can substantially improve discourse parsing.
-They not only encode large amounts of lexical and conceptual information, addressing problem (ii) but their increased context windows permits an incremental parsing strategy to address problem (iii).
-Such a strategy makes sense cognitively and linguistically, as prior discourse structure clearly has an effect on inferences about which subsequent discourse units are attached (the link task) and how they are attached—i.e.
-what semantic relations link them (the link+relation task) Asher and Lascarides, (2003).
-Further, an incremental approach naturally applies to downstream applications like conversational agent, in which the information about the previously inferred discourse structure must be updated and available in real time.
-We develop an incremental parsing model LLaMIPa (LLaMa Incremental Parser), by fine-tuning LLaMA3, that uses previously inferred discourse structure to simultaneously predict the link and the link+relation tasks between new discourse units.
-We show LLaMIPa substantially outperforms the state of the art on both tasks.
+They not only encode large amounts of lexical and conceptual information, addressing problem (ii), but their increased context windows permit an incremental parsing strategy to address problem (iii).
+Such a strategy makes sense cognitively and linguistically, as prior discourse structure has an effect on inferences about where subsequent discourse units are attached (the link task), and how they are attached, i.e., which types of semantic relations link them (the link+relation task) Asher and Lascarides, (2003).
+Further, an incremental approach naturally applies to downstream applications such as conversational agents, in which the information about the previously inferred discourse structure must be available and updated in real time.
+We develop an incremental parsing model by fine-tuning Llama3 Dubey et al., (2024).
+Llamipa (Llama Incremental Parser) uses previously inferred discourse structure to simultaneously predict the links and relation types between new discourse units.
+We show that Llamipa substantially outperforms the state-of-the-art on both the link and link+relation tasks.
 We demonstrate Llamipa’s performance using three datasets annotated in the style of SDRT (Segmented Discourse Representation Theory Asher,, 1993; Asher and Lascarides,, 2003).
-Our central dataset is the Minecraft Structured Dialogue Corpus (MSDC; Thompson et al.,, 2024), a corpus of situated, chat-based, task-oriented dialogues, but we show the generality of our approach by separately fine-tuning on STAC, a corpus of situated multi-party chats from an online game Asher et al., (2016).
-We further test (STAC-based) LLaMIPa on an out of domain dataset, Molweni Li et al., (2020).
-Not only does LLaMIPa outperform prior models, but it handles particularly complex structural features of our situated, multiparty datasets, addressing problem (i).
-In particular, LLaMIPa can identify cases in which a single EDU has two parent EDUs—a situation that arises naturally in both multiparty and situated dialogue where a speaker can respond to multiple channels (conversational or multimodal) at once.
+Our central dataset is the Minecraft Structured Dialogue Corpus (MSDC; Thompson et al.,, 2024), a corpus of situated, chat-based, task-oriented dialogues, but we also show the generality of our approach by separately fine-tuning on STAC, a corpus of situated multiparty chats from an online game Asher et al., (2016).
+We further test Llamipa fine-tuned on STAC on an out of domain dataset, Molweni Li et al., (2020).
+Not only does Llamipa outperform previous discourse parsing models, but it handles particularly complex structural features of our situated, multiparty datasets, addressing problem (i).
+In particular, Llamipa can identify cases in which a single EDU has two parent EDUs, a situation that arises naturally in both multiparty and situated dialogue, where a speaker can respond to multiple channels (conversational or multimodal) at once.
 Most current SDRT based parsers cannot handle such structures Shi and Huang, (2019); Liu and Chen, (2021); Wang et al., (2021); Chi and Rudnicky, (2022).
@@
 Sections 5 and 6 describe our model and results.
-Section 7 describes ablation experiments, and then we conclude in Section 8.
+Section 7 describes our ablation experiments, and we conclude in Section 8.
 2 Background
-Discourse representations for a text or conversation often take the form of tree or graph structures that are built recursively from EDUs, where EDUs are clauses or subclausal units that serve as minimal, linguistic constituents in discourse relations Marcu, (1999).
-Such relations, such as Explanation, Question-Answer Pair or Acknowledgement, are represented as linking EDUs in the discourse graph.
+Discourse representations for a text or conversation often take the form of tree or graph structures that are built recursively from discourse units or EDUs, where EDUs are clauses or subclausal units that serve as minimal, linguistic constituents in discourse relations Marcu, (1999).
+These relations, such as Explanation, Question-Answer Pair or Acknowledgement, are represented as linking EDUs in a discourse graph.
 There are two main theories that have investigated complete discourse structures for texts: Rhetorical Structure Theory (RST; Mann and Thompson,, 1987) and SDRT (Asher,, 1993; Asher and Lascarides,, 2003).
-RST builds discourse trees over EDUs with the semantic relation labels on the head nodes dominating subtrees.
-SDRT provides a dependency style, weakly connected directed acyclic graph (DAG) structure in which typed relations are represented as labeled arcs in the graph, and the terms of the relations are EDUs or elementary event units (EEUs).
+RST assigns relations between contiguous EDUs which are connected hierarchically to further EDUs in a treelike structure.
+SDRT provides a dependency style, weakly connected directed acyclic graph (DAG) in which typed relations are represented as labeled arcs in the graph, and the terms of the relations are EDUs or elementary event units (EEUs).
 EEUs represent salient aspects of the nonlinguistic context in which a conversation takes place, and may be connected to the discourse DAG.
 SDRT also allows subgraphs, or complex discourse units (CDUs) to act as terms of semantic relations.
-Discourse parsing requires addressing three tasks.
-The segmentation task involves decomposing a text and (descriptions of) actions into EDUs and EEUs.
+SDRT is the only framework to date that handles multiparty conversation.
+Whatever the underlying paradigm, discourse parsing requires addressing three tasks.
+First, the segmentation task, which involves decomposing a text and (descriptions of) actions into EDUs and EEUs.
 The link task involves identifying pairs of EDUs and/or EEUs that are semantically related via discourse relations—that is, the arcs in the discourse graph.
-The link+relation task involves identifying the labels for each attached pair.
+Finally, the link+relation task involves identifying the labels for each attached pair.
 3 Related work
-Traditionally with SDRT style parsers, gold segmentation is assumed Muller et al., (2012); Afantenos et al., (2015); Perret et al., (2016); Shi and Huang, (2019); Liu and Chen, (2021); Chi and Rudnicky, (2022), and parsing follows a pipeline architecture in which discourse units are first linked and then labels are subsequently predicted for these links.
-While we assume gold segmentation, our proposed architecture combines the attachment and labeling tasks.
-Several strategies have evolved to take account of the complex information required for discourse parsing within a dependency based paradigm like that proposed by SDRT or DLTag (Forbes et al.,, 2003).
+Traditionally, SDRT style parsers assume gold discourse unit segmentation Muller et al., (2012); Afantenos et al., (2015); Perret et al., (2016); Shi and Huang, (2019); Liu and Chen, (2021); Chi and Rudnicky, (2022), and parsing follows a pipeline architecture in which discourse units are first linked, and then labels are predicted for those links.
+Our proposed architecture assumes gold discourse unit segmentation, but combines the attachment and labeling tasks.
+Several strategies have evolved to harness the complex information required for discourse parsing within the dependency based paradigm proposed by SDRT or DLTag (Forbes et al.,, 2003).
 Research has shown that semantic features of EDUs are crucial for solving the link+relation task Mann and Thompson, (1987); Asher and Lascarides, (2003).
-Older approaches to multiparty conversation Afantenos et al., (2015); Perret et al., (2016) thus used manual features about EDU pairs and simple ML models to build a local model to predict link and link+relation label in a pipeline fashion.
-A local model uses information just about pairs of EDUs that are to be attached; it does not exploit information about the discourse structure around that pair.
-Afantenos et al., (2015); Perret et al., (2016) also added decoding mechanisms with post hoc constraints on the output structures, following Muller et al., (2012).
-Although exploiting little to no information about other components of the discourse structure, such a model can still be competitive Bennis et al., (2023).
+Thus, older approaches to multiparty conversation Afantenos et al., (2015); Perret et al., (2016) input manually designed features of the EDU pairs to simple ML models to predict link and link+relation label successively.
+These local models use information local to the pairs of EDUs that are to be attached, but do not exploit information about the surrounding discourse structure.
+In order to leverage some more global information, Afantenos et al., (2015); Perret et al., (2016) apply special decoding mechanisms with post hoc constraints on the output EDU pairs, following Muller et al., (2012).
+Although local models exploit little to no information about other components of the discourse structure, such a model can still be competitive Bennis et al., (2023).
 We do a detailed comparison between an encoder based local parsing model and Llamipa.
-Shi and Huang, (2019) were the first to obtain significant discourse parsing results using a deep learning architecture on STAC.
+Shi and Huang, (2019) were the first to obtain competitive discourse parsing results using a deep learning architecture on STAC.
 They attempted to capture incremental and contextual effects in their model by training a supplemental Structured Encoder to incrementally update attachment paths (sequences of parent-child EDUs).
-However, Wang et al., (2021) showed that the model obtains similar scores with or without this encoder.
+However, Wang et al., (2021) show that the model obtains similar scores with or without this encoder.
 Liu and Chen, (2021) develop an approach that uses a pre-trained RoBERTa model Liu et al., (2019) to provide embeddings for EDUs.
@@
 Their model uses two linear layers for link and relation prediction.
-Wang et al., (2021) provide the Structure Self-Aware Graph Neural Network (SSA-GNN), which is a complex GNN-based architecture and model that uses both EDU and edge embeddings.
+Wang et al., (2021) provide the Structure Self-Aware Graph Neural Network (SSA-GNN), which is a complex GNN-based architecture and model that uses both EDU and relation embeddings.
 The model uses a Hierachichal GRU gate for contextual EDU representations.
@@
 None of the models by Shi and Huang, (2019); Liu and Chen, (2021); Guz et al., (2020); Wang et al., (2021); Chi and Rudnicky, (2022) predict multiple parents for an EDU (or EEU).
-The only previously developed parser that predicts multiple parents during attachment is Bennis et al., (2023); Thompson et al., (2024)’s model BERTLine.
+The only previously developed parser that predicts multiple parents during attachment is Bennis et al., (2023)’s model BERTLine.
 BERTLine uses BERT embeddings for EDUs and EEUs, along with a linear layer for features like speaker change and distance information, to build a local model and predicts link and link+relation in pipeline fashion.
@@
 Thompson et al., (2024) uses BERTLine on the MSDC, which makes it a natural baseline for our work on situated, multiparty conversation.
-Recent work has investigated using LLMs to predict discourse structure, by finetuning LLaMA2 7b, 13b, and 70b for RST-style parsing Maekawa et al., (2024).
-The reported results on RST corpora are not directly comparable to ours, but they are significantly lower numerically.
-Chan et al., (2024); Yung et al., (2024); Fan et al., (2024); Henning and Stede, (2024) have investigated discourse parsing through prompting with large GPT-style models to produce PDTB (Prasad et al.,, 2008) or other discourse annotations.
-The results are significantly worse than dedicated discourse parsers; testing with 3-shot learning on a version of the STAC corpus (STAC-L; see Section 4), Chan et al., (2024) reports an F1 score of 21.3 for link prediction and 7.4 for relation prediction by ChatGPT.
+Recent work investigates using LLMs to predict discourse structure by finetuning Llama2 7b, 13b, and 70b for RST-style parsing Maekawa et al., (2024).
+The reported results show improvements over the state-of-the-art for RST corpora analogous to those we find using Llamipa on SDRT corpora.
+Chan et al., (2024); Yung et al., (2024); Fan et al., (2024) have investigated discourse parsing through prompting with large GPT-style models to produce PDTB (Prasad et al.,, 2008) or other discourse annotations.
+These models fare worse than dedicated discourse parsers: Chan et al., (2024) test with 3-shot learning on a version of the STAC corpus (see Section 4 for more corpora details) and report a near 50 point drop in F1 for link and link+relation prediction compared to Shi and Huang, (2019).
 4 Data sets
-MSDC.
-The MSDC provides full discourse structures in the style of SDRT to dialogues in the Minecraft Dialogue Corpus (MDC; Narayan-Chen et al.,, 2019).
-In the MDC, dialogues involve a builder, whose task is to build block formations in a simplified Minecraft world, and an architect, who guides the builder by giving instructions through chat.
-The MSDC annotates discourse relations between both chat moves (EDUs) and linguistic descriptions of the builder’s pick and place actions (EEUs), provided by the MDC logs.
-The MSDC’s multimodal nature presents many opportunities to simultaneously respond to a conversational move and a building move, and indeed, the MSDC has numerous discourse units with multiple parents (MPDUs).
-Figure 1 shows an excerpt from the MSDC with two examples.
-For instance, after placing two blocks, the builder asks “like that?”.
-The architect’s reply, “not quite,” not only serves to answer the builder’s question but simultaneously to correct the builder’s last moves.
-The following two elaborating discourse moves fill out the full content of the Correction.
-Figure 1 shows another important aspect of the MSDC, namely that builders tend to perform sequences of pick and place moves uninterrupted by chat moves.
-These action sequences are grouped together as CDUs in the MSDC, indicated by dashed boxes in Figure 1.
-Bennis et al., (2023), Thompson et al., (2024) merged these highly repetitive sequences of nonlinguistic moves into a single EEU to prevent them from drowning out the more complex and interesting linguistic interactions in the corpus.
+MSDC. The MSDC provides full discourse structures in the style of SDRT for dialogues in the Minecraft Dialogue Corpus (MDC; Narayan-Chen et al.,, 2019).
+The MSDC represents conversational moves as well as builder move sequences, individuated into block placements and removals.
+Figure 1 shows an excerpt from an MSDC dialogue containing a few examples of discourse units with multiple parents, or MPDUs.
+In one such example, after placing two blue blocks, the builder asks, “like that?”.
+The architect’s reply, “not quite,” not only serves to answer the builder’s question, but also corrects the builder’s last moves.
+The EDU for “like that?” is an MPDU that is semantically linked to both linguistic and nonlinguistic units.
+Such MPDUs are an important feature of the MSDC, as they facilitate the interplay of multimodal contributions in a single discourse structure.
+Another noteworthy feature of the MSDC is how nonlinguistic actions are annotated.
+Builder moves are usually prompted by architect instructions, but often transpire in longer sequences of block placements and removals, which sometimes overlap with other linguistic chat moves.
+These action sequences are grouped together in single complex discourse units (CDUs) in the MSDC, indicated by the dashed boxes in Figure 1.
+Bennis et al., (2023), Thompson et al., (2024) merged these highly repetitive sequences of nonlinguistic moves into a single EEU to prevent the repetitive connections between them from drowning out the more complex and interesting linguistic interactions in the corpus.
 We used the MSDC with this modification.
@@
 | EEU | 0/0 | 11566/1156 | 4687/1473 |
-| MDPU | 838//86 | 999/96 | 4789/1476 |
-| MPDU-3 | 94/9 | 95/6 | 85/35 |
+| MPDU | 838//86 | 999/96 | 4789/1476 |
+| MPDU=3 | 94/9 | 95/6 | 85/35 |
 | MPDU>3 | 11/0 | 7/0 | 1/3 |
@@
 STAC has two versions: STAC-L, in which only relations between chat moves (EDUs) are annotated in the absence of the context provided by nonlinguistic events, and a situated version, STAC-Sit, in which both chat moves and (descriptions of) game moves (EEUs) are annotated in the presence of an integrated linguistic and nonlinguistic context.
-STAC contains MPDUs due to both multimodal interactions and multiparty discourse, in which one player can simultneously respond to multiple players at once.
+Like MSDC, STAC contains MPDUs due to multimodal interactions, but also features MPDUs which arise from multiparty interaction, in which one player can simultaneously respond to multiple players at once.
 Relevant statistics for STAC and the MSDC are found in Table 1.
-STAC-Sit, like the MSDC, also contains sequences of game-only moves.
-Bennis et al., (2023) compressed such sequences into individual EEUs and we follow this choice.
+STAC-Sit, like the MSDC, also contains sequences of moves containing only EEUs.
+Bennis et al., (2023) compressed these sequences into individual EEUs and we follow this choice.
 Unlike the MSDC, however, STAC contains numerous CDUs for chat moves and for chat/game interactions (as opposed to CDUs of action sequences only).
-Following Bennis et al., (2023) and most of the literature on SDRT parsing, we did not treat CDUs directly but used the “flattening" procedure outlined in Muller et al., (2012) that converts a structure with CDUs into one with only EDUs and EEUs.
-A final simplication of the original STAC corpus involved removing EEUs that neither interact with EDUs n or figure in EEU sequences that interact with EDUs.
-This choice, like the one to compress EEU sequences, was made to avoid drowning out the more complex and interesting data involving linguistic EDUs and their interactions with EEUs.
+Following Bennis et al., (2023) and most of the literature on SDRT parsing, we do not treat CDUs directly, but use the “flattening” procedure outlined in Muller et al., (2012) that converts a structure with CDUs into one with only EDUs and EEUs.
+A final simplification of the original STAC corpus involved removing EEUs that neither interact with EDUs nor figure in EEU sequences that interact with EDUs.
+This choice, like the one to compress EEU sequences, is made to avoid drowning out the more complex and interesting data involving linguistic EDUs and their interactions with EEUs.
 Molweni.
-The Molweni corpus contains SDRT graphs for multiparty interactions in the Ubuntu Chat Corpus Lowe et al., (2015) a collection of exchanges from an online forum for discussing issues related to Ubuntu.
+The Molweni corpus contains SDRT graphs for multiparty interactions in the Ubuntu Chat Corpus Lowe et al., (2015), a collection of exchanges from an online forum for discussing issues related to Ubuntu.
 Unlike the MSDC and STAC, Molweni does not contain CDUs, EEUs, or MPDUs.
-We used the version of Molweni Test re-annotated by Li et al., (2024).
+We use the Molweni-clean test set, a version of Molweni re-annotated by Li et al., (2024).
 5 Llamipa: An Incremental Parser
-LLaMIPa leverages the sequence to sequence architecture and the enhanced contextual reasoning capacities of generative LLMs by introducing an incremental discourse parsing strategy.
-As it generates a discourse graph for a conversation , the LLM should base its link and relation predictions for an EDU in on the graph it has already constructed for EDUs preceding .
-Formally, let be a conversation consisting of a sequence of EDUs or EEUs.
-Our LLM takes as input a set of of discourse and (possibly) event units and a discourse graph .
-At step 0, is , the head EDU, and , the null discourse graph.
-After that, is a set , where 1 and where is the window size, set to ; and contains links for .
-At each step, the model predicts a graph that extends by integrating link+relation predictions for all of the EDUs in the next dialogue turn or, if the next move is an action sequence, the EEU representing this sequence.
-This process repeats until LLaMIPa arrives at the final unit of , effectively creating a “sliding context window” that will eventually slide over an entire dialogue, permitting the model to potentially exploit a large amount of contextual information.
-We represent relation instances as EDU index pairs connected by a relation type, e.g.
-RES(0,1), for the instance of a Result relation holding between EDUs 0 and 1.
-During training, each subsequent instance in a dialogue adds to the context the previous turn EDUs and EEUs, but we take the gold, rather than predicted, discourse structure between those units to make our input graph .
-is specified as a list of the relations between these units in the same format as the predicted relations.
-We used a sliding context widow of 15 EDUs to optimize the number of training samples while capturing the most number of discourse links that have statistical support.
-As discourse links follow an inverse logarithmic curve Afantenos et al., (2015), there is little statistical support for links of distance longer than 15 in the SDRT corpora.
-Using our corpora formatted for incremental relation prediction, we finetuned LLaMA-2 Touvron et al., (2023), and LLaMA-3222https://llama.meta.com/llama3/ to produce LLaMIPa2 and LLaMIPa3.
-In order to do this without training the full model weights, we use QLoRA Dettmers et al., (2024) to train for 3 epochs.
+Llamipa leverages the enhanced contextual reasoning capacities of generative LLMs by introducing an incremental discourse parsing strategy.
+As it generates a discourse graph for a conversation, the LLM bases its link and relation predictions for an elementary unit based on the graph it has already constructed for units preceding .
+Formally, let be a sequence where denotes an EDU or EEU.
+Our LLM takes as input a set , containing EDUs and EEUs, and a discourse graph .
+At step , is , the head EDU, and , the null discourse graph.
+After that, is a set , where and where is the window size, and contains links connecting .
+At each step, the model extends the previous discourse graph by integrating link and relation predictions for all of the EDUs or EEUs in the current dialogue turn.
+In the data, we represent relation instances as elementary unit index pairs connected by a relation type, e.g.
+RES(0,1), for the instance of a Result relation holding between units and .
+During training, each sample contains the previous EDUs and EEUs within the window size, the gold structure, which is a list of relation instances between these elementary units, and finally, the units of the current turn.
+The model is trained to predict the gold relation instances that connect the units of the current turn with previous ones.
+As LLMs can handle longer input sequences, we use a window size which allows the model to exploit a large amount of contextual information.
+Since discourse links follow an inverse logarithmic curve Afantenos et al., (2015), there is little statistical support for links of distance longer than in the SDRT corpora.
+We finetune Llama 2 Touvron et al., (2023), and Llama 3 Dubey et al., (2024) to produce Llamipa2 and Llamipa3 using QLoRA Dettmers et al., (2023).
+Both the models are trained for 3 epochs.
 The model parameter details are in Table 6 in Appendix A.
-The finetuned models will be made publicly available.
-Once LLaMIPa is finetuned on the training corpus, the generation procedure uses predicted discourse structure instead of the gold structures during inference (See Figure 2).
-Proceeding turn by turn, LLaMIPa produces a complete discourse structure for a dialogue in which both links and relations are predicted.
-The LLaMIPa3 model and generation code is available here333https://huggingface.co/Kqte/LLaMIPa.
+We release Llamipa3 publicly through the Hugging Face platform222https://huggingface.co/linagora/Llamipa.
+We test our models in two settings.
+In the first setting, the models predict the relation instances for samples containing the gold discourse structure.
+In the second, the model uses its predicted relation instances instead of the gold structure in the input sample.
+That is, at each step, the structure fed to the model is just the accumulated relation instances predicted within the window up until that step (see Figure 2).
+This incremental setting provides a clear idea of the model’s ability to build a complete dialogue structure from scratch.
 6 Results
-We first ran our finetuning experiments on the MSDC corpus, and assessed LLaMIPa’s performance on link and relation type prediction when we include the gold dialogue structure in each sample context.
-We then tested LLaMIPa by using the previously predicted relations in the context instead of the gold ones, as in the generation step described above, in order to have our incremental parser build a predicted structure from beginning to end.
-Table 2 shows that LLaMIPa3 significantly outperforms BERTLine and outperforms LLaMIPa2 both using gold structure (LLaMIPa3+g vs.
-LLaMIPa2+g) and predicted structure (LLaMIPa3+p vs.
-LLaMIPa2+p).
-Below we use the better performing LLaMIPa3+p for other evaluations.
-To test the generalizability of the model, we also finetuned LLaMIPa3 on STAC-Sit and compared it to BERTLine on MSDC and STAC-Sit.
-For both link prediction and link relation prediction, LLaMIPa shows a gain of almost 10 F1 points or 11% on link and 10.5 F1 points or 15% on link relation labeling.
-LLaMIPa posted even more impressive gains for predictions of certain relations.
-When we look at F1 breakdown by relation type in Table 5, the largest relative increase comes from improved predictions for the Narration and Correction relation types, which Thompson et al., (2024) cite as the most difficult to predict given their particular characteristics.
-We then compared our model to a broader range of SDRT discourse parsers on the STAC linguistic version (STAC-L), as previous work focused on this version of the corpus.
-Table 4 shows that LLaMIPa3 finetuned on STAC-L and evaluated on STAC-L Test scored significantly higher on link prediction and got a statistically significantly higher score on link + relation prediction over the state-of-the-art parser Chi and Rudnicky, (2022).
-Nevertheless, LLaMIPa3 has substantially lower performance on STAC-L (linguistic) than on either of the situated corpora.
-This mirrors human efforts at annotation.
-Asher et al., (2016, 2021) report that annotating the STAC linguistic data set was more difficult; annotators felt that information needed to determine the discourse structure was missing.
-We also investigated LLaMIPa3 for transfer learning, finetuning on STAC-L and testing on the Molweni Cleaned Test corpus of Li et al., (2024).
-While superior to other models, in comparison to Li et al., (2024), LLaMIPa3 fared worse on link prediction but was far superior with respect to link + relation prediction.
-Overall, LLaMIPa has several advantages.
-It simplifies previous attach-then-type pipeline methods by predicting link and relation labels simultaneously.
-Previous research has given theoretical arguments that co-predicting attachment and relation labels should give better results Asher and Lascarides, (2003), as well as empirical evidence for this claim Bennis et al., (2023).
-Our model also outperforms prior models on all relevant corpora and pretty much all tasks we studied.
-We hypothesize that this performance is due to the richer contextual information supplied to LLaMIPa, a hypothesis we now explore further.
+We run finetuning experiments using the MSDC corpus, and assess Llamipa’s performance on link and link+relation prediction under the two test settings described in Section 5.
+We consider only the well-formed outputs during evaluation, but note that Llamipa outputs only well-formed relation instances for all experiments.
 | MSDC | STAC Situated | |||
@@
 | BERTLine | 0.7870 | 0.6901 | 0.7667 | 0.6788 |
-| Distance | 2 | 3 | 4 | 5 | 6 | 7 | 8 | 9 | 10 | 11 | 12 | 13 | 14 | 15 |
-| LLaMIPa-3 | 0.90 | 0.88 | 0.82 | 0.76 | 0.82 | 0.80 | 0.79 | 0.71 | 0.90 | 0.76 | 0.50 | 0.40 | 1.0 | n/a |
-| BERTLine | 0.70 | 0.61 | 0.43 | 0.32 | 0.38 | 0.02 | 0.21 | 0.28 | 0.08 | n/a | n/a | n/a | n/a | n/a |
 | STAC-L | STAC- L/Molweni-clean | |||
@@
 | Llamipa-3+p | 0.775 | 0.607 | 0.712 | 0.405 |
-| BERTLine | LLaMA3-Local | LLaMIPa3+p | LLaMIPa3+g | LLaMIPa-rand | No Structure | |
+Table 2 shows that Llamipa3 significantly outperforms BERTLine and outperforms Llamipa2 both using gold structure (Llamipa3+g vs. Llamipa2+g) and predicted structure (Llamipa3+p vs. Llamipa2+p).
+To test the generalizability of our approach, we also finetune Llamipa on STAC-Sit and compare the Llamipa3+p outputs to BERTLine.
+Llamipa3+p shows a gain of F1 points on link prediction and F1 points on link+relation prediction.
+As mentioned in Section 5, the predicted context setting best demonstrates the desired functionality of an incremental parser.
+Since Llamipa3+p give the best performance in this setting, we use it for all evaluations and henceforward refer to it using the shorthand Llamipa.
+Performance on STAC-L.
+In order to compare Llamipa to a broader range of SDRT discourse parsers, we tested it on STAC-L, as prior work has only used this linguistic version of STAC.
+Table 3 shows that Llamipa finetuned on STAC-L and evaluated on the STAC-L test set scores substantially higher on both link and link+relation prediction than the state-of-the-art parser Chi and Rudnicky, (2022).
+We note that Llamipa has a lower performance on STAC-L than on either of the situated corpora.
+This reflects the findings of Asher et al., (2016, 2021), who reported that annotating the STAC linguistic data set was more difficult; annotators felt that the nonlinguistic information needed to determine the discourse structure was missing (see Section 4).
+| Distance | 2 | 3 | 4 | 5 | 6 | 7 | 8 | 9 | 10 | 11 | 12 | 13 | 14 | 15 |
+| Llamipa-3 | 0.90 | 0.88 | 0.82 | 0.76 | 0.82 | 0.80 | 0.79 | 0.71 | 0.90 | 0.76 | 0.50 | 0.40 | 1.0 | n/a |
+| BERTLine | 0.70 | 0.61 | 0.43 | 0.32 | 0.38 | 0.02 | 0.21 | 0.28 | 0.08 | n/a | n/a | n/a | n/a | n/a |
+We also test Llamipa on transfer learning.
+Following Li et al., (2024), we finetune on STAC-L and test on their Molweni-clean test set.
+While Llamipa outperforms all models on STAC-L, it fares worse on link prediction than Li et al., (2024) on Molweni, though it does better on link+relation prediction; a counterintuitive result given that link+relation prediction is the harder task.
+Taking a closer look at the link prediction results in Li et al., (2023), we see their model performs well on links of distance , but drops after that ( accuracy for distance , and for distances greater than ).
+Given that Molweni is a relatively simple corpus where most of the links are adjacent, their model excels on the link task.
+But for relation type prediction, Li et al., (2024) use the traditional method, which assigns relation types after establishing the attachment structure, and using only local information.
+Llamipa surpasses their model on relation type prediction, which shows that an approach predicting link and relation simultaneously performs better than a local pipeline approach, even for shallower discourse structures.
+Performance on MSDC.
+Llamipa shows impressive gains for predictions of certain relation types.
+Looking at F1 breakdown by relation type in Table 5, we see the largest relative increase comes from improved predictions for Narration and Correction, which Thompson et al., (2024) cite as the most difficult to predict given that they are heavily context dependent.
+Narrations are frequently quite long (with a significant proportion above distance ), and difficult to capture by local parsers like BERTLine Thompson et al., (2024).
+Table 4 shows Narration F1 at distances between and EDUs.
+Like other local parsers, BERTLine requires a distance cutoff that balances ignoring longer relations with mitigating the class imbalance caused by numerous unlinked candidates; for SDRT corpora, this is distance .
+As indicated in Section 5, Llamipa continues to make good predictions up to distance .
+Further, Llamipa is able to predict many instances of the Correction relation, which also has a number of long distance instances in the MSDC.
+Table 5 shows that its performance is F1 points, which is 42 points higher than and more than double that of BERTLine.
+We hypothesize that Llamipa has a high F1 on Narration and Correction because it has access to rich semantic information about the discourse context: it can use the previous discourse moves as well as an explicit representation of their discourse structure.
+This hypothesis coheres with our understanding of how Narration functions in the MSDC.
+Narrations connect high-level instructions between which participants clarify, question and elaborate on those instructions, as well as correct missteps (see Figure 1).
+While there is some statistical similarity in their surface forms, the important cues for isolating a new instruction, and hence a Narration attachment, involve certain patterns of Result and Acknowledgement links in the prior context, as well as the presence of (nonlinguistic) block placements.
+This pattern is regular enough that one can find most Narrations programmatically as with Thompson et al., (2024)’s “Second Pass”.
+This takes predictions from a first pass of the local model for Results and Acknowledgements, then applies an algorithmic approach to predict long distance Narration links.
+We tried this second pass for Llamipa, however there was no improvement on F1 for Narration scores, indicating that Llamipa has already internalized this pattern during training.
+Llamipa offers improvements over BERTLine for many other relation types which are typically not considered long distance.
+Table 5 shows Llamipa has higher scores for Result, Acknowledgment, Elaboration, Question-answer Pair, Clarification Question, and Comment.
+We hypothesize that predictions on these also profit from complex contextual information, something that we discuss in more detail in Section 7.
+Conversely, for relations where Llamipa gives slightly poorer results—Contrast, Alternation, and Question-Elaboration—we surmise that these types can be inferred reliably just from the local information given by the related EDUs or EEUs.
+Overall, Llamipa has several advantages.
+It simplifies previous attach-then-type pipeline methods by predicting link and relation labels simultaneously.
+Previous research has given theoretical arguments that co-predicting attachment and relation labels should give better results Asher and Lascarides, (2003), as well as empirical evidence for this claim Bennis et al., (2023).
+And as we have demonstrated, Llamipa outperforms other parsers on all relevant corpora.
+| BERTLine | Llama3-local | Llamipa3+p | Llamipa3+g | Llamipa+rand | Llamipa+ | |
 | Result | 0.85 | 0.86 | 0.90 | 0.91 | 0.83 | 0.72 |
@@
 7 The role of context
-A particularly remarkable demonstration of LLaMIPa’s discourse parsing prowess is its success in predicting Narration instances in the MSDC.
-Narrations are frequently quite long (with a significant proportion above distance 10), and so difficult to capture by local parsers like BERTLine Thompson et al., (2024).
-Table 3 shows Narration F1 at distances between 3 and 14 EDUs.
-Like other local parsers, BERTLine requires a distance cutoff that balances ignoring longer relations with mitigating the class imbalance caused by numerous unlinked candidates; for SDRT corpora, this is distance 10.
-As explained in Section 5, LLaMIPa allows for training up to distance 15, and impressively, it continues to make (somewhat steadily good) predictions up to distance 15.
-LLaMIPa also has very strong performance in predicting instances of the Correction relation, which also has a significant number of long distance instances in the MSDC.
-Table 5 shows that its performance is more than double BERTLine’s on this relation; its F1 score for Correction is over 42 F1 points more than BERTLine’s.
-We hypothesize that LLaMIPa has a high F1 on Narration and Correction because it has access to rich semantic information about the discourse context.
-It can use the meaning of previous discourse moves as well as an explicit representation of their discourse structure.
-This hypothesis coheres with our understanding of how Narration functions in the MSDC.
-Narrations connect high-level instructions between which participants clarify, question and elaborate on those instructions, as well as correct missteps (see Figure 1).
-While there is some statistical similarity in their surface forms, the important cues for isolating a new instruction, and hence a Narration attachment, involve certain patterns of Result and Acknowledgement links in the prior context, as well as the presence of (nonlinguistic) block placements.
-This pattern is regular enough that one can find some Narrations algorithmically as in Thompson et al., (2024)’s “Second Pass”: taking predictions from a first pass of the local model, Thompson et al., then apply an algorithmic approach to predict long distance Narration links from the local model’s predictions.
-We also tried this second pass for LLaMIPa, however there was no improvement on F1 for Narration scores, which shows that LLaMIPa had already learned this rule during training.
-Besides its performance on Narration and Correction, LLaMIPa offers significant improvements over BERTLine for many other relation types, even those whose instances are typically not very long distance.
-Table 5 shows LLaMIPa has significantly higher scores for Result, Acknowledgment, Elaboration, Question-answer Pair, Clarification Question and Comment.
-We hypothesize that predictions on these too profit from complex contextual information.
-LLaMIPa produces poorer results for relations such as Contrast, Alternation and Question-Elaboration.
-This may be because instances of these relations can be inferred reliably just from information given by the related EDUs or EEUs.
+So far, the hypothesis has been that Llamipa’s state-of-the-art performance, especially for longer distance relations, is due to its ability to process richer contextual information.
+Whether this is a function of the superior embeddings conferred by the underlying LLM, or a function of a larger context window that allows a longer dialogue history and explicit discourse structure as prediction input, is something we explore further in this section.
 First ablation: embeddings vs.
-context.
-While the previous discussion suggests that LLaMA’s larger context window is responsible for its better performance, we wanted to check this hypothesis further.
-So we finetuned LLaMA3-8b over pairs of EDUs as input to predict links and relation types between pairs of EDUs.
-This produces a local model roughly comparable to BERTLine’s, though BERTLine’s architecture is different.
-Table 5 shows that, compared to BERTLine, LLaMA3-Local gave stronger performance without using the prior discourse context for link and relation prediction.
-However, when compared to LLaMIPa3, LLaMA3-Local has significantly lower link and link+relation labeling predictions overall.
-LLaMIPa3’s relation predictions for many relations are also significantly better—indeed sometimes vastly superior as in the F1 for Narration.
-This supports our conclusion that although LLaMA’s superior embeddings play a role, many relations, and especially relations like Narration, require more contextual information like that available to LLaMIPa.
-Further ablations: removing structure.
-Given the result of our first ablation experiment, we wanted to investigate further whether LLaMIPa was using the discourse structure for its predictions, or was solely relying on the longer input sequence of EDUs and EEUs.
-Our set up for LLaMIPa makes the assumption that it is the prior discourse structure that is relevant.
-In Table 5, there is a small but consistent drop between gold and predicted contexts in F1 performance for all relation types.
-This supports the conclusion that LLaMIPa takes the discourse structure of its context into account and that the quality of that structure affects prediction.
-To provide more support for this conclusion, we performed another ablation in which we perturbed the structure at each increment by randomizing both the relation type and EDU index for each relation.
-We then tested LLaMIPa (finetuned in the usual way) on this input.
-LLaMIPa-rand showed a significant drop in performance from LLaMA3-Local for link and across almost all relation types (with the exception of Contrast and Question Elaboration).
-The drop for Narration and Correct is sharp, as well as for Question-answer Pair.
-We then performed a variant ablation, in which we removed all structure from LLaMIPa’s input context, leaving only a sequence of EDUs and EEUs.
-The performance of LLaMIPa on this data set (in the column “No Structure” of Table 5) shows another significant drop both in link and in link+relation predictions.
-The drop in relation labeling in addition holds for almost all relations, and a dramatic drop, especially for Narration, Correction, Clarification Question and Question-Answer Pair.
-This provides strong evidence that LLaMIPa needs to the use discourse structure and not just the EDU/EEU input sequence of the context to optimize its predictions for links and relation labels and that discourse structure is crucial for predicting some relations.
+context size.
+This ablation tests whether the LLM embeddings for EDUs and EEUs account for Llamipa’s performance.
+We finetune Llama3-8b over pairs of elementary units as input to predict links and relation types, resulting in a local model roughly comparable to BERTLine.
+Table 5 shows that, compared to BERTLine, Llama3-local still performed better without using the prior discourse context for link and link+relation prediction.
+However, when compared to Llamipa, Llama3-local still has considerably lower predictions overall.
+Llamipa’s relation predictions for many relations are also better, notably for Narration.
+This supports our conclusion that although Llama’s superior embeddings play a role, many relations, and especially relations like Narration, require the information available in Llamipa’s larger context window.
+Second ablation: perturbing input structure.
+Even if Llamipa requires the input available in a larger context window, it is still an open question whether it uses the discourse structure provided in this window.
+We already have some indication that Llamipa does use the structure when we look at Llamipa’s performance given a gold discourse structure for a context versus its predicted structure, as described in Section 5.
+The results of this comparison are shown in Table 5: there is a small but consistent drop between gold and predicted structure contexts in F1 performance for all relation types.
+To investigate this further, we perform a second ablation where we perturb the input structure for each test sample by randomizing both the relation type and EDU index for each relation instance in the structure.
+When we run Llamipa on this randomized input (denoted as Llamipa+rand in Table 5) we see a drop in performance from Llamipa3+p for link and link+relation for most relation types, with a substantial drop for Narration, Correction, and Question-answer Pair.
+Third ablation: removing input structure.
+In this ablation, we removed all structure from Llamipa’s input context, leaving only a sequence of elementary units.
+The performance of Llamipa on this data set (in the column Llamipa+ of Table 5) shows an even larger drop in both link and link+relation predictions compared to Llamipa+rand.
+Similar to Llamipa+rand, we see a substantial drop for the relation types Narration, Correction, and Question-Answer Pair.
+Additionally, the performance dropped drastically for Clarification Questions.
+Fourth ablation: targeted perturbations.
+The results of the previous ablations support the hypothesis that the structure provided in the context window does have an effect on relation instance predictions.
+This final ablation investigates whether Llamipa predicts context dependent relations as well as it does because it leverages (or “reasons” about) the relation instances in the context structure in a manner consistent with our semantic intuitions.
+Here, we focus on instances of Question-answer pair (QAP) and Correction.
+QAP attachments provide a straightforward example of context dependence.
+In the MSDC, it is common that if a prediction QAP(y,z) is made for a dialogue turn z then CLARIFQ(x,y), CONFQ(x,y), or QELAB(y,z) is in its context.
+That is, the questions that are answered in the QAP relations tend to be questions asked to clarify, confirm or elaborate upon the preceding context.
+To test whether Llamipa’s predictions conform to this pattern, we isolate the test samples where Llamipa correctly predicts QAP (Table 5), and remove all question relation types from the structure given in the context of these samples.
+We re-run the generation step on this amended sample set, and see a point drop in F1, from to , for QAP predictions, and no change for all other relation types (see Figure 3).
+Correction relations are among the most difficult to identify in the MSDC, as discussed in Section 6.
+Although the language used for Corrections usually features some form of negation—No, Not like that, Oops!—-deciding where it attaches requires making inferences about how the content of the corrective move applies to a more global information state.
+Although this is beyond the capability of a discourse parser like Llamipa, there are still structural regularities in the MSDC that can help with retrieving Corrections.
+For instance, most Corrections come about when the Builder makes the wrong moves in response to an instruction, and the Architect then gives a further instruction, which results in moves that address this by fixing the last wrong moves.
+In the discourse structure, these interactions create recurring patterns, “correction triangles” (Figure 4).
+These triangles provide a way to check Llamipa’s sensitivity to CORR(x,y) type relations in the context structure in cases where it must predict RES(y,z) CORR(x,z), the last two relations in a triangle.
+As with the Question-answer Pair test, we single out each test sample in which Llamipa correctly predicts the second CORR; but instead of removing CORR(x,y), we change it to ACK(x,y) (Acknowledgement(x,y)).
+This leaves the structure intact, but gives it a contrary sense.
+We re-run the generation step and see a point drop in F1, from to , for Correction predictions, and no change for all other relation types.
 8 Conclusions
-In this paper, we presented LLaMIPa3, an incremental SDRT style discourse parser that uses its prior discourse structure predictions to predict links and relation labels for new discourse units.
-We finetuned both LLaMa2-13b and LLaMA3-8b models to produce LLaMIPa2 and LLaMIPa3.
-We showed that LLaMIPa3 performs better than other parsers that use SDRT data sets.
-It excels in particular on data sets that feature conversation that contains both linguistic moves and nonlinguistic contexts that help interpret them.
-We also compared LLaMIPa3 with a local model discourse parser based on LLaMA3-8b as well as on variants of LLaMIPa3 that had differing levels of access to the discourse context.
-LLaMIPa3 had significantly higher F1 scores for attachment and relation labeling on the data sets we used for evaluation than all of these.
-The differential performance of LLaMIPa3 on STAC-L versus STAC-Sit, which reflects impressions of human annotators, suggests that nonlinguistic contexts are important for stable interpretation of situated conversation.
-We humans typically invent a nonlinguistic context for a conversation when one isn’t given.
-But such an invention is subjective and may contribute to variations in understanding and prediction.
-We suspect something like this might be true for LLMs too.
-In future work, we hope to use LLaMIPa3 to help conversational robots to improve their performance in interactive tasks.
+Llamipa is an incremental SDRT style discourse parser that uses its prior discourse structure predictions to predict links and relation labels for the discourse units in new conversation turns.
+As such, it can be naturally assimilated into downstream applications like conversational agents.
+After finetuning both Llama2-13b and Llama3-8b models to produce Llamipa2 and Llamipa3, we show that Llamipa3 has substantially higher F1 scores than other parsers that use SDRT data sets on the link and link+relation tasks.
+We confirm that its performance is due to access to larger swaths of context, which are not available to parsers that use a local model.
+We further find evidence to support the hypothesis that Llamipa uses the explicit representations of structure in its context to predict new relation instances.
+Notably, Llamipa excels on data sets that feature conversation containing both linguistic and nonlinguistic moves (actions, in the MSDC) that help interpret them.
+This indicates that a parser like Llamipa can furnish important discourse information to builder action models; preliminary demonstrations of this can be found in Chaturvedi et al., (2024).
 Limitations
@@
 Llamipa, like previous parsers, presumes that the EDUs/EEUs are already segmented.
-However extant work on automatic segmentation suggests that this task is relatively tractable on diverse data sets Muller et al., (2019) including even corpora of spoken conversation (Gravellier et al.,, 2021; Prevot et al.,, 2023)
-Though ablations show that LLaMIPa uses discourse structure, it remains unclear how it leverages that structure to predict relations like Narration, Corrections, and QAP.
+However, existing work on automatic segmentation suggests that this task is relatively tractable on diverse data sets Muller et al., (2019) including even corpora of spoken conversation (Gravellier et al.,, 2021; Prevot et al.,, 2023)
 Like all supervised models, Llamipa is limited to making predictions on data sets that are similar to its training data.
@@
 Our work here has been to improve the capacities of AI models to understand the semantic structure of discourse and conversation.
-We see no direct ethical concerns that arise from this work, though we hope that a better grasp of the semantic structure of conversation may enable models to eventually produce less pragmatically grounded or conversational hallucinations.
+We see no direct ethical concerns that arise from this work, though we hope that a better grasp of the semantic structure of conversation may enable the output of models to be more pragmatically grounded.
+Acknowledgements
+For financial support, we thank the National Interdisciplinary Artificial Intelligence Institute ANITI (Artificial and Natural Intelligence Toulouse Institute), funded by the French ‘Investing for the Future–PIA3’ program under the Grant agreement ANR-19-PI3A-000.
+This project has also been funded by the France 2030 program and is funded by the European Union - Next Generation EU as part of the France Relance.
+We also thank the projects COCOBOTS (ANR-21-FAI2-0005) and DISCUTER (ANR-21-ASIA-0005), and the COCOPIL “Graine” project of the Région Occitanie of France.
+This work was granted access to the HPC resources of CALMIP supercomputing center under the allocation 2016-P23060.
 Appendix A Appendix: Model Parameters
@@
 | quantization for Llama3 | 4-bit NormalFloat |
-Table 6 gives the hyperparameters used for finetuning LLaMIPa-2 and LLaMIPa-3 along with the computing resources.
-We adapted the finetuning code from the following repository444https://github.com/mlabonne/llm-course/blob/main/Fine_tune_Llama_2_in_Google_Colab.ipynb.
-We provide LLaMIPa3 trained on MSDC as well as the MSDC test data here555https://huggingface.co/Kqte/LLaMIPa
+Table 6 gives the hyperparameters we use for finetuning Llamipa-2 and Llamipa-3 along with the computing resources.
+We adapt the finetuning code from the following repository333https://github.com/mlabonne/llm-course/blob/main/Fine_tune_Llama_2_in_Google_Colab.ipynb.
 Appendix B Appendix: STAC Linguistic data
-We constructed STAC-L, our STAC linguistic corpus, from the STAC data available on the STAC website (https://https://www.irit.fr/STAC/corpus.html) and adopted standard procedures for flattening CDUs.
-However, we found that our data set did not quite match that of Bennis et al., (2023) in terms of numbers of EDUs.
+We derive STAC-L from the STAC data available on the website (https://https://www.irit.fr/STAC/corpus.html) by adopting standard procedures for flattening CDUs.
+However, we find that our data set did not quite match that of Bennis et al., (2023) in terms of numbers of EDUs.
 Shi and Huang, (2019) also report a different number of EDUs from what we have and from what Bennis et al., (2023) report.
-When we finetuned and tested LLaMIPa on the STAC linguistic only data available in the BERTLine repository 666https://github.com/zineb198/LineBert, we had lower scores than what we report in Table 4.
+When we finetune and test Llamipa on the STAC-L data available in the BERTLine repository 444https://github.com/zineb198/LineBert , we had lower scores than what we report in Table 3.
 Llamipa-3+p had an F1 for link of 75.4 and a link+relation F1 of 57.7 on this version of STAC linguistic.
-On the transfer learning from STAC linguistic to Molweni Cleaned Test, LLaMIPa-3+p had an F1 of 68.5 for link and 38.7 for link+relation.
+On the transfer learning from STAC linguistic to the Molweni-clean test set, Llamipa-3+p had an F1 of 68.5 for link and 38.7 for link+relation.
+We make our version of STAC-L available on the STAC corpus website cited above.